OpenStudy (anonymous):

Suppose a parabola has an axis of symmetry at x = –7 , a maximum height of 4 and also passes through the point (–6, 0). Write the equation of the parabola in vertex form.

OpenStudy (p0sitr0n):

y =a(x-h)^2+k y= a(x+7)^2+4 0 = a(-6+7)^2+4 -4=a y=-4(x-h)^2+4

OpenStudy (anonymous):

Thank you . Is that the answer ? (:

OpenStudy (p0sitr0n):

ya, Your welcome

OpenStudy (anonymous):

@andreascott yes its correct answer